OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 2 points) Convert 3750 mL to the unit dm3. 3.75 x 104 dm3 3.75 dm3 3.75 x 10-2 dm3 3.75 x 106 dm3 -------------------------------------------------------------------------------- Question 2 (Multiple Choice Worth 2 points) Convert 1.03 g/mL to the unit g/m3. 1.03 x 106 g/m3 1.03 x 102 g/m3 1.03 x 10-2 g/m3 1.03 x 10-6 g/m3 --------------------------------------------------------------------------------

OpenStudy (fifciol):

question 1: 1L=1dm^3 1mL= 10^-3 dm^3 so 3750 mL =3,75 dm^3

OpenStudy (fifciol):

question 2: 1mL =10^-3 dm^3 =10^-3 (0,1m)^3=10^-3 *10^-3 m^3=10^-6 m^6 so 1.03 g/mL = 1.03 x 106 g/m3